OpenStudy (anonymous):

Taylor has a credit card with an interest rate of 24.5% compounded monthly that she must pay off in seven years. If her balance on the credit card is $2,345, what must her monthly payment be?

OpenStudy (tkhunny):

I = 0.245 -- Annual Percentage Rate j = I/12 = 0.020416667 - Monthly Percentage Rate v = 1/(1+j) = 0.979991833 - Monthly Discount Rate \(2345.00 = Pmt\cdot(v + v^{2} + ... + v^{84}) = Pmt\cdot\dfrac{v - v^{85}}{1-v}\)

OpenStudy (tkhunny):

You will struggle until you can build these formulas on the fly!